# Designing a quasi-experiment to study the clinical impact of adaptive risk prediction models

### v1

@@ Abstract
Clinical risk prediction is a valuable tool for guiding healthcare interventions toward those most likely to benefit. Yet, evaluating the pairing of a risk prediction model with an intervention using randomized controlled trials (RCTs) presents substantial challenges to today’s healthcare systems. This makes quasi-experimental designs, which can offer nearly the same level of evidence as an RCT, an attractive alternative. However, existing quasi-experimental designs do not allow models and thresholds to adapt. As a result, they struggle to serve new populations, meet emerging trends, and address practical issues. To address this gap, we introduce regression discontinuity designs for evaluating risk prediction models paired with specific interventions in a modern healthcare system. In our designs, treatment is assigned when predicted risk crosses a defined threshold, with the design explicitly accommodating adaptations in both the model and threshold. We account for the interference that arises from these adaptations to estimate the local average treatment effect in a valid and efficient way. To that end, we characterize interference and provide sufficient conditions for identification. Estimators are introduced, and their performance is evaluated in a simulation that emulates how cardiovascular risk calculators could guide interventions in primary care settings.

### v2

@@ Abstract
Clinical risk prediction is a valuable tool for guiding healthcare interventions toward those most likely to benefit. Yet, evaluating the pairing of a risk prediction model with an intervention using randomized controlled trials presents substantial challenges, making quasi-experimental designs an attractive alternatives. Existing designs, however, assume that both the model and the decision rules used to trigger interventions (typically a risk threshold) remain fixed. This limits their utility in modern healthcare, where both are routinely updated. We introduce a regression discontinuity framework that accommodates adaptation in both the model and the risk threshold. We precisely characterize the form of interference introduced by these adaptations and exploit this structure to establish conditions for identification and thus design estimation strategies. The key idea is to define counterfactual risks—the scores patients would have received under hypothetical reorderings—thereby restoring local exchangeability and enabling valid estimation of the local average treatment effect. Our estimator leverages the fact that, although counterfactual risk vectors grow with time, they typically lie in a low-dimensional space. In simulations of cardiovascular prevention programs, we show that our method accurately recovers treatment effects even as thresholds adapt to meet operational or clinical targets and models are updated to align predicted and observed outcomes or to exclude demographic predictors such as race.